OpenStudy (anonymous):

Solve. 6 – 4x = 4(3 – x) – 6 (Points : 4) all real numbers 3 5 8

OpenStudy (studygurl14):

6 - 4x = 4(3 - x) -6 First, distribute the 4 6 - 4x = 12 - 4x - 6 Do 12 - 6 = 6 6 - 4x = 6 - 4x So...all real numbers